A hospital needs 20L of 5% solution. They have 20% solution and 4% solution available. How many L of each 20% and 4% solutions would they have to mix in order to get 20L of 5% solution.
Here's how I've done so far:
x+y=20L
.2x
.04y
.05 (x=y)
.2x+.04y=.05 (x+y)
(I'm not sure if I'm even going the right direction here and this is where I stop knowing what to do)

Answer by mananth(16946)

percent ---------------- quantity
Solution I 20.00% ---------------- x liters
Solution II 4.00% ------ 20 - x liters
Mixture 5.00% ---------------- 20
20
20.00% x + 4.00% ( 20 - x ) = 5.00% * 20
20 x + 4 ( 20 - x ) = 100
20 x + 80 - 4 x = 100
20 x - 4 x = 100 - -80
16 x = 20
/ 16
x = 1.25 liters 20.00% Solution
18.75 liters 4.00% solution

m.ananth@hotmail.ca